OpenStudy (anonymous):

Which point lies on the graph of this function? y=-x^2+4x-2 (-1,3) (0,2) (1,1) (2,-2) Can someone help me with this I don't know where to start.

OpenStudy (campbell_st):

just substitute the x value for each point into the equation and see if you get the y value from the point

OpenStudy (anonymous):

Ohh okay thank you

OpenStudy (anonymous):

none of them worked

OpenStudy (anonymous):

wait I messed up I forgot the negative sign out in front

OpenStudy (campbell_st):

I think its the negative to start check (1, 1) y = -(1)^2 + 4(1) - 2 so y = -1 + 4 - 2 y = 1 so it looks like (1, 1) is the answer